# A probe of the maximum energetics of fast radio bursts through a prolific repeating source

### v1

@@ Acknowledgements
[MHz] [MHz] subband [MHz] [Jy] threshold [Jy ms] threshold [Jy ms] threshold [$\mathrm{10^{30}\,erg\,Hz^{-1}}$] [hr]
@@
g Hours spend on source between MJD 59867 and 59983 (15 October 2022 and 08 February 2023).
@@
Table 2: Best Schechter fit results from different studies
This work $41.32^{+0.45}_{-0.30}$ $-1.12^{+0.17}_{-0.13}$ 24.4
Ryder et al. (2023)21,c $41.7^{+0.2}_{-0.2}$
Shin et al. (2023)16 $41.38^{+0.51}_{-0.50}$ $-1.30^{+0.7}_{-0.4}$ 5.0
James et al. (2022)15 $41.26^{+0.72}_{-0.22}$ $-1.95^{+0.18}_{-0.15}$ 0.5/4.4/21.981
Luo et al. (2020)20,d $42.08^{+0.30}_{-0.06}$ $-1.79^{+0.31}_{-0.35}$ 0.55
aEnergy assuming a conical adopted bandwidth of $1$ GHz.
bBest fitted slope parameters on a differential distribution.
cBased on the sample of James et al. (2022)15.
eFluence limits on the different observational campaigns.
Figure 1: Overview of the FRB 20220912A observing campaign
Each coloured block represents an observation with a telescope at a certain frequency. The top two panels show $22$ days and the bottom two panels show $36$ days with the associated MJD and and calendar dates on the x-axis. The broken y-axis show the frequency range. The vertical dotted lines denote detections of bursts where black illustrates detections at L-band ($1.4~{}\mathrm{GHz}$) and red detections at P-band ($0.3~{}\mathrm{GHz}$). The black bar in the top panels indicate the overlapping observing window with FAST and NRT.
Figure 2: Ratio of fluences computed from SFXC- and digifil-generated filterbanks.
The fluence for each burst was measured three different ways due to digitisation artefacts present in the data. These filterbanks were made for bursts detected with Westerbork and Toruń using SFXC, digifil and digifil with a scattered power correction (SPC) applied. Here we plot the ratio of these measured fluences for L-band (left panel) and P-band (right panel). For L-band we find that due to digitisation effects the fluence of the burst can be underestimated for up to $40\%$ (SFXC) and overestimated for $10\%$ (digifil). Applying the SPC algorithm to correct for digitisation effects is therefore essential. Even though we detect bright ($>1000$ Jy ms) bursts at P-band, these digitisation effects are less apparent, this most likely due to longer dispersive sweep of the burst and the low sensitivity of the Westerbork P-band receiver.
Figure 3: DM determination for bursts detected at P-band.
Peak signal-to-noise (SN) versus dispersion measure (DM) curve for our brightest detection at P-band, burst B128-Wb. When applying the DM used to correct for dispersive delay for bursts detected at L-band, $219.37$ pc cm-3, there still was a residual sweep present for bursts detected at P-band, as illustrated in the middle panel. By correcting the burst for range of DM values and fitting a Gaussian to the SN-curve we find a best fit DM of $219.375$ pc cm-3 and shown in the right panel.
Figure 4: Cumulative burst energy distribution for bursts detected by NRT.
Left: The cumulative energy distribution for NRT, also shown in Figure 2, fitted with two power laws. The vertical dotted black line is the estimated completeness threshold and the vertical dotted red line indicates the turnover point as estimated using the powerlaw package. The solid black line corresponds to the determined breakpoint of the distribution at $E_{\rm{break}}=3.2~{}\times~{}10^{30}$ erg/Hz. Right: the power law index ($\gamma_{\rm{C}}$) as a function of the spectral energy density calculated by a maximum-likelihood estimation. The horizontal dotted black lines indicate the slopes of the power law used in the left panel.
Figure 5: Predicted optical afterglows for maser shock model
@@
Table 1: Schechter function parameter results
Results and best fit parameters from the MCMC analysis on fitting a Schechter function to the differential burst distribution by varying the fluence/energy threshold for bursts detected between MJD 59869 and 59910. The coloured green row denotes the the completeness threshold and parameter values which were used in Figure 3.
$\sigma$ Threshold fluence Threshold spectral energy Threshold Total Bursts $N_{\textrm{Wb}}$ $N_{\textrm{St}}$ $N_{\textrm{Tr}}$ $N_{\textrm{ATA}}$ $N_{\textrm{Nc}}$ $N_{\textrm{Fast}}$ N $\gamma_{D}$ $E^{\textrm{char}}$ $\chi^{2}_{\nu}$
[Jy ms] [$\mathrm{10^{30}\,erg\,Hz^{-1}}$] [$\mathrm{10^{32}\,erg\,Hz^{-1}}$]
10 $16.10$ $2.18$ 146 25 68 4 13 32 4 $9.68^{+24.87}_{-8.30}$ $-1.14^{+0.14}_{-0.10}$ $2.54^{+6.87}_{-1.33}$ $1.13$
14.55a $23.81$ $3.22$ 130 23 64 2 9 30 2 $12.44^{+28.52}_{-10.14}$ $-1.13^{+0.16}_{-0.12}$ $2.16^{+4.22}_{-1.06}$ $0.85$
15b $24.40$ $3.30$ 125 22 62 2 8 29 2 $13.55^{+32.61}_{-10.93}$ $-1.12^{+0.17}_{-0.13}$ $2.09^{+3.78}_{-1.04}$ $0.78$
20 $32.45$ $4.40$ 112 20 55 2 7 27 1 $20.32^{+40.40}_{-16.08}$ $-1.05^{+0.20}_{-0.15}$ $1.68^{+2.70}_{-0.76}$ $0.66$
30 $48.80$ $6.61$ 93 12 47 1 7 26 0 $40.06^{+66.02}_{-31.28}$ $-0.90^{+0.28}_{-0.23}$ $1.19^{+1.68}_{-0.49}$ $0.50$
Table 2: Burst properties
@@
The complete table and an additional table with properties per component for each burst can be found in the supplementary material.
@@
[MJD] [Jy ms] of components [ms] [$\mathrm{10^{29}\,erg\,Hz^{-1}}$] [$\mathrm{10^{32}\,erg\,s^{-1}\,Hz^{-1}}$] [MHz] [MHz]

### v2

@@ 2 Observations
[MHz] [MHz] subband [MHz] [Jy] threshold [Jy ms] threshold [Jy ms] threshold $\mathrm{10^{30}\,erg\,Hz^{-1}}$] [hr]
@@
g Hours spend on source between MJD 59867 and 59983 (2022 October 15 and 2023 February 08).
@@
2.1 Westerbork, Onsala & Toruń
@@
We used Heimdall to search for bursts in the filterbank files. In order to minimise the amount of false positive candidates we set a signal-to-noise (S/N) threshold of $7$ and limit the dispersion measure (DM) search to within ± $\pm\,50$ pc cm-3 of the reported value of $\mathrm{DM_{FRB}=220}$ pc cm-3 (Mckinven & CHIME/FRB Collaboration, 2022). We mitigate radio frequency interference (RFI) by applying a static mask, which excises certain frequency channels known to contain RFI. The identified burst candidates are then classified using the machine learning classifier FETCH (Agarwal et al., 2020), where we make use of models A &\& H and set a detection threshold of $50\,\%$. The burst candidates that have a reported probability of at least $50\,\%$ in one of the two models are then all manually inspected. As a fail-safe we also manually inspect all burst candidates that have a reported DM within $5$ pc cm-3 of the expected DM.
Using the radiometer equation we calculate both the detection and completeness thresholds for the observations (Cordes & McLaughlin, 2003). The detection threshold represents the minimal fluence that our telescopes are sensitive to, while the completeness threshold represents the minimal fluence where we expect to detect (almost) all bursts. For the detection threshold we assume a 7-$\sigma$ detection with a canonical adopted FRB width of $1~\textrm{ms}$. For the completeness threshold we assume a conservative 15-$\sigma$ threshold and take a burst width of $3~\textrm{ms}$ since we know that bursts originating from FRB 20220912A have generically longer durations than one millisecond (Konijn et al., 2024). The various detection and completeness thresholds of the instruments, per observing band, are listed in Table 1.
2.2 Stockert
[…]
2.3 Observational pointing
[…]
3 Analysis and Results
We detected 130 bursts from FRB 20220912A. Of these, 114 unique bursts were detected at $1.4\,\mathrm{GHz}$, including 16 bursts that were detected by multiple telescopes simultaneously. Remarkably, the highest-energy bursts from FRB 20220912A contributed to $22.0^{+15.6}_{-10.3}\,\%$ of the all-sky rate of FRBs at L-band ($\mathcal{F}>500\mathrm{~Jy~ms}$). We also detected $16$ unique bursts at $330\,\mathrm{MHz}$, while no bursts were detected at $4.7\,\mathrm{GHz}$. Although we sometimes had simultaneous coverage between all observing bands, see Figure 7, we did not detect any burst at multiple frequency bands. Each burst is labeled with an ID, $\rm{B}x$, numbered in order of their arrival time, and followed by the telescope code to indicate which instrument detected the burst (e.g., B15-Tr). A subset of bursts, preferentially those with high S/N, is shown in Figure 1; a complete overview of all dynamic spectra is provided in the Supplementary material.
Figure 1: Dynamic spectra, time and temporal profiles for a subset of bursts. Each subfigure consists of three panels. Shown in the top panel is the burst-id, the time- and frequency resolution at which the data is plotted and the time profile of the burst. The colored bars represent the width for each component of a burst whereas the color of the bars correspond the instrument used to detect the burst. Purple corresponds to Stockert (St), orange to Toruń (Tr) and blue to Westerbork (Wb). The side panel shows the temporal profile which is the sum over the time axis, but only under the colored bars. The white vertical lines are masked channels at the edges of the subbands or the presence of radio frequency interference (RFI) which are indicated by red ticks. The bursts have been corrected for dispersive effects where we used a value of $219.37$ pc cm-3 for bursts detected at $1.4~\mathrm{GHz}$ (L-band) and $219.73$ pc cm-3 for bursts detected at $0.3~\mathrm{GHz}$ (P-band). For Stockert this correction was applied incoherently (between frequency channels) and for Toruń and Westerbork this correction was applied incoherently and coherently (within frequency channels). The dynamic spectra of all detected bursts are available as part of the Supplementary material.
3.1 Burst properties
To correct for the dispersive delay we used a single DM for all bursts detected at L-band ($1.4\,\mathrm{GHz}$). The value we used is $219.375$ pc cm-3, which was determined by Hewitt et al. (2023) from analysis of the extremely bright and broad-band microshots in a burst that was co-detected by NRT (B2 in that paper) and Westerbork (B52-Wb here). While the DM of FRB 20220912A is known to vary (Hewitt et al., 2023), we find that such variations have a negligible effect on the inferred fluences and hence energetics of the bursts, which is the main focus of this paper. Hence, we do not attempt to derive an optimal DM for each burst, but see §B.1.
The burst properties we use in our analyses were determined using the filterbank files created with digifil and the SPC-algorithm, previously referred to as Method III. RFI was mitigated by manually masking affected frequency channels for all bursts using the tools psrzap and pazi from the DSPSR software package. Additionally, we also zapped the edges of the subbands because of the drop in sensitivity at these frequencies (see, e.g., B35-st and B15-tr in Fig. 1).
To measure the time and frequency extent of the bursts we follow our earlier work (Kirsten et al., 2024). We manually determined the start and stop times of each individual component of a burst. For these time ranges we calculate the 2D auto-correlation function in time and frequency. We then fit a 1D Gaussian to the spectra and time series and determine the width both in time and frequency. We define the burst frequency extent to be the full extent of the observing frequency if the full width at half maximum (FWHM) of the Gaussian fit is more than $75\,\%$ of the total bandwidth. The fluence of a burst is determined using the radiometer equation and calculating the flux density per time bin while summing over the on-time region for each component (Cordes & McLaughlin, 2003). An overview of burst properties can be found in Table 2.
Additionally, we also measured the fluences using filterbank files created by SFXC (Method I) and only digifil (Method II) in order to compare and quantify the digitisation effects that are described in more detail in §B.2. The ratios of fluences between the three methods are shown in Figure 9. We find that we encounter saturation effects that underestimate the fluences by up to $40\,\%$ (SFXC) and overestimate them by $10\,\%$ (digifil) for bursts detected at L-band (left panel). For bursts detected at P-band (right panel) the saturation effect is less apparent even though we detect bright bursts (>1000>1000 Jy ms). This lack of saturation is most likely due to the longer dispersive sweep of the burst and the relatively low sensitivity of the Westerbork P-band receiver.
@@ 3.1 Burst properties
Table 2: Burst properties. The complete table and an additional table with properties per component for each burst can be found in the supplementary material.
@@
[MJD] [Jy ms] of components [ms] $\mathrm{10^{29}\,erg\,Hz^{-1}}$] $\mathrm{10^{32}\,erg\,s^{-1}\,Hz^{-1}}$] [MHz] [MHz]